# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v1

@@ 2.4 Posterior Inference
Here, for a better illustration of the proposed framework, we rewrite the prior for $\bm{\theta}_{jl}$ as the generative model with observation vector $\bm{0}$. We note that since the distribution $f\left(\bm{\theta}\mid\bm{\Omega}\right)$ is the indicator function, there is no need to estimate $\bm{\theta}$ separately. Figure 2 gives a graphical representation of the Bayesian model.
Figure 2: A graphical representation of the proposed approach.
Based on (2), we develop an effective Expectation Maximization (EM) algorithm for posterior inference of the parameter set $\bm{\Psi}=\operatorname{vec}\left(\bm{\beta},\bm{\tau},{\bm{\mu}},{\bm{\Omega}}\right)$. First, we introduce an indicator $\rho_{k,i}=I\left\{g_{i}=k\right\}$ and take $\bm{\gamma}$ and $\bm{\rho}=\left(\rho_{1,1},\cdots,\rho_{K,n}\right)^{\top}$ as latent variables. Then, for initialization, we set $\bm{\beta}=\bm{0}$ and $\bm{\Omega}_{k}=\bm{I}$ for $k=1,\ldots,K$, where $\bm{I}$ is an identity matrix, randomly partition the subjects into $K$ subgroups, and initialize ${\tau}_{k}$ and $\bm{\mu}_{k}$ as the precision parameter of residuals and the vector of predictor centers in the $k$th subgroup, respectively. To obtain $\bm{L}^{(jl)}$ for each $j<l$, we adopt the initial values of the precision matrices, choose $\epsilon=0.001$, and update $\bm{L}^{(jl)}$ accordingly based on $\omega_{k,jl}$’s in the following iterations.
In the E step, given the estimations $\left({\bm{\Psi}}^{(m-1)},{\bm{\pi}}_{\bm{g}}^{(m-1)}\right)$ in the $(m-1)$th step, compute the conditional expectations of $\bm{\gamma}$ and $\bm{\rho}$ as $\gamma_{k,jl}^{(m)}=\left(\frac{p_{1}}{2v_{1}}\exp\left\{-\frac{\left|\omega_{k,jl}^{(m-1)}\right|}{v_{1}}\right\}+\frac{1-p_{1}}{2v_{0}}\exp\left\{-\frac{\left|\omega_{k,jl}^{(m-1)}\right|}{v_{0}}\right\}\right)^{-1}$ $\frac{p_{1}}{2v_{1}}\exp\left\{-\frac{\left|\omega_{k,jl}^{(m-1)}\right|}{v_{1}}\right\}$ and ${\rho}_{k,i}^{{(m)}}=\frac{{\pi}_{k}^{(m-1)}f_{k}\left(t_{i},\bm{x}_{i},\delta_{i}\mid\bm{\Psi}_{k}^{(m-1)}\right)}{\sum_{k^{\prime}=1}^{K}{\pi}_{k^{\prime}}^{(m-1)}f_{k}\left(t_{i},\bm{x}_{i},\delta_{i}\mid\bm{\Psi}_{k^{\prime}}^{(m-1)}\right)}$, where $\bm{\Psi}_{k}=\operatorname{vec}\left({\bm{\mu}}_{k},{\bm{\Omega}}_{k},{\bm{\beta}}_{k},{\beta}_{0k},{{\tau}}_{k}\right)$.
Then the conditional expectation of the log-posterior can be calculated as:
$L_{n}\left({\bm{\Psi}},\bm{\pi}_{\bm{g}}\mid{\bm{\Psi}}^{(m-1)},\bm{\pi}_{\bm{g}}^{(m-1)}\right)$ $=$ $E_{\bm{\gamma},\bm{\rho}\mid{\bm{\Psi}}^{(m-1)},\bm{\pi}_{\bm{g}}^{(m-1)}}\log f\left(\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\gamma},\bm{g}\mid\bm{t},\bm{\delta},\textbf{X},\bm{c};\bm{\tau},\bm{\pi_{g}}\right)$ (3)
$\equiv$ $L_{1,n}\left(\bm{\beta},\bm{\tau}\right)+L_{2,n}\left(\bm{\mu},\bm{\Omega}\right)+L_{3,n}\left(\bm{\pi}_{g}\right)+\text{Const},$
with $L_{1,n}\left(\bm{\beta},\bm{\tau}\right)=\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{K}{{\rho}_{k,i}^{{(m)}}}\left[(1-\delta_{i})\log S_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k};\beta_{0k},{\tau}_{k}\right)+\delta_{i}\log f_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k};\beta_{0k},{\tau}_{k}\right)\right]-\lambda_{1}\sum\limits_{k=1}^{K}\sum\limits_{j=1}^{p}|\beta_{k,j}|$, $L_{2,n}\left(\bm{\mu},\bm{\Omega}\right)=\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{K}\frac{{\rho}_{k,i}^{{(m)}}}{2}\left[\log\left|\bm{\Omega}_{k}\right|-\left(\bm{x}_{i}-{\bm{\mu}}_{k}\right)^{\top}\bm{\Omega}_{k}\left(\bm{x}_{i}-{\bm{\mu}}_{k}\right)\right]-\lambda_{2}\sum\limits_{k=1}^{K}\sum\limits_{j=1}^{p}|\mu_{k,j}|-\sum_{j<l}\left\{\sum\limits_{k=1}^{K}\left[\frac{\gamma^{(m)}_{k,jl}}{v_{1}}+\frac{1-\gamma^{(m)}_{k,jl}}{v_{0}}\right]\left|\omega_{k,jl}\right|+\frac{u\bm{\omega}_{jl}^{\top}\bm{L}^{(jl)}\bm{\omega}_{jl}}{2}\right\}$, and $L_{3,n}\left(\bm{\pi}_{\bm{g}}\right)=\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{K}{\rho}_{k,i}^{{(m)}}\log\pi_{k}$.
In the M step, we update the parameters $\left(\bm{\Psi},\bm{\pi}_{\bm{g}}\right)$ by maximizing $L_{1,n},L_{2,n}$, and $L_{3,n}$, separately.
Denote $u_{k,i}^{(m-1)}={f_{Z}\left(t_{i}\mid\bm{x}_{i},{\bm{\beta}}_{k}^{(m-1)};{\beta}_{0k}^{(m-1)},{\tau}_{k}^{(m-1)}\right)}/{S_{Z}\left(t_{i}\mid\bm{x}_{i},{\bm{\beta}}_{k}^{(m-1)};{\beta}_{0k}^{(m-1)},{\tau}_{k}^{(m-1)}\right)}$ and $\tilde{z}_{i,k}^{(m-1)}=\delta_{i}t_{i}+(1-\delta_{i})\left[\bm{x}_{i}^{\top}{\bm{\beta}}_{k}^{(m-1)}/{\tau}_{k}^{(m-1)}+{\beta}_{0k}^{(m-1)}/{\tau}_{k}^{(m-1)}+{u_{k,i}^{(m-1)}}/{\left({\tau}_{k}^{(m-1)}\right)^{2}}\right]$. The optimization for $\bm{\tau}$ and $\bm{\beta}$ proceeds as follows (and we refer to Supplementary Materials for more details).
  1. (1)
    Compute $\tau_{k}^{(m)}=\frac{\widetilde{b}_{k}^{(m)}+\sqrt{\left(\widetilde{b}_{k}^{(m)}\right)^{2}+4\widetilde{a}_{k}^{(m)}\widetilde{c}_{k}^{(m)}}}{2\widetilde{a}_{k}^{(m)}}$, where $\widetilde{b}_{k}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\tilde{z}_{i,k}^{(m-1)}\left[(1-\delta_{i})(u_{k,i}^{(m-1)}-\right.$ $\left.\tau_{k}^{(m-1)}t_{i}+r_{ik}^{(m-1)})+r_{ik}^{(m-1)}\right]$, $\widetilde{c}_{k}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left[\delta_{i}+(1-\delta_{i})\left(u_{k,i}^{(m-1)}-\tau_{k}^{(m-1)}t_{i}+r_{ik}^{(m-1)}\right)\right]$ with $r_{ik}^{(m-1)}=\bm{x}_{i}^{\top}\bm{\beta}_{k}^{(m-1)}+\beta_{0k}^{(m-1)}$, and $\widetilde{a}_{k}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left(\tilde{z}_{i,k}^{(m-1)}\right)^{2}$.
  2. (2)
    Conduct a coordinate descent (CD) algorithm to optimize $L_{1,n}$ with respect to $\bm{\beta}_{k}$. Specifically, for $j=1,\cdots,p$, carry out the following steps sequentially. Compute $h_{kj}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}x_{i,j}^{2}$ and $res_{-kj}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left(\tau_{k}^{(m)}\tilde{z}_{i,k}^{(m-1)}-r_{ik}^{(m)}\right)x_{i,j}+h_{kj}^{(m)}\beta_{k,j}^{(m-1)}$, where $r_{ik}^{(m)}=r_{ik}^{(m-1)}+x_{i,j}\beta_{k,j}^{(m)}-x_{i,j}\beta_{k,j}^{(m-1)}$. Then update $\beta_{k,j}^{(m)}=\frac{res_{-kj}^{(m)}-\lambda_{1}\operatorname{sgn}\left(res_{-kj}^{(m)}\right)}{h_{kj}^{(m)}}$ if $\left|res_{-kj}^{(m)}\right|>\lambda_{1}$ and $\beta_{k,j}^{(m)}=0$ otherwise.
  3. (3)
    Compute ${\beta}^{(m)}_{0k}=\frac{\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left({\tau}_{k}^{(m)}z_{i,k}^{(m-1)}-\bm{x}_{i}^{\top}{\bm{\beta}}_{k}^{(m)}\right)}{\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}}.$
Then, we maximize $L_{2,n}$ with respect to $\bm{\mu}$ and $\bm{\Omega}$ as follows.
  1. (1)
    For each $k=1,\ldots,K$, compute ${\bm{\mu}}_{k}^{(m)}=\begin{cases}\frac{grad_{k,j}^{(m)}-\lambda_{2}\operatorname{sgn}\left(grad_{k,j}^{(m)}\right)}{\omega_{k,jj}^{(m-1)}\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}},&\text{if }~{}\left|grad_{k,j}^{(m)}\right|>\lambda_{2},\\ 0,&\text{otherwise, }\end{cases}$ where $grad_{k,j}^{(m)}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left[\sum_{l\neq j}\left(x_{i,l}-\mu_{k,l}^{(m-1)}\right)\omega_{k,lj}^{(m-1)}+x_{i,j}\omega_{k,jj}^{(m-1)}\right].$
  2. (2)
    For $\bm{\Omega}$, we have $\left({\bm{\Omega}}_{1}^{(m)},\ldots,{\bm{\Omega}}_{K}^{(m)}\right)=\arg\max_{\bm{\Omega}_{1},\ldots,\bm{\Omega}_{K}}\left\{\sum_{k=1}^{K}{n_{k}}\left[\log\left|\bm{\Omega}_{k}\right|-\operatorname{tr}\left(\widetilde{\bm{S}}_{k}\bm{\Omega}_{k}\right)\right]\right.$ $\left.-\sum_{j<l}\left[2\sum_{k=1}^{K}\left(\frac{\gamma^{(m)}_{k,jl}}{v_{1}}+\frac{1-\gamma^{(m)}_{k,jl}}{v_{0}}\right)\left|\omega_{k,jl}\right|+u{\bm{\omega}_{jl}^{\top}\bm{L}^{(jl)}\bm{\omega}_{jl}}\right]\right\},$ where $n_{k}=\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}$ and $\widetilde{\bm{S}}_{k}={\sum_{i=1}^{n}{\rho}_{k,i}^{{(m)}}\left(\bm{x}_{i}-\bm{\mu}_{k}^{(m)}\right)\left(\bm{x}_{i}-\bm{\mu}_{k}^{(m)}\right)^{\top}}/n_{k}$ is a pseudo sample covariance matrix. This optimization problem can be efficiently solved using the ADMM algorithm. We refer to Supplementary Materials for details.
Final, maximizing $L_{3,n}$ with respect to $\bm{\pi}_{g}$ subject to $\sum_{k=1}^{K}\pi_{k}=1$ yields $\pi_{k}^{(m)}=\frac{\sum_{i}{\rho}_{k,i}^{{(m)}}}{n}.$
Through the iterative E and M steps, convergence is typically attained within a reasonable number of iterations. To maintain consistency in model selection, with the estimate $\hat{\bm{\Omega}}_{k}$ obtained by the EM algorithm, we further adopt a threshold strategy and have the final estimate $\widetilde{\omega}_{k,jl}=\hat{\omega}_{k,jl}{I}\left\{\left|\hat{\omega}_{k,jl}\right|>0.001\right\}$, following the literature [13, 11]. We refer to Theorem 1 for more details. To improve stability, we explore multiple random initializations of subjects’ subgroup memberships, and the final estimates are selected based on the smallest Akaike Information Criterion (AIC), as defined below.
For the proposed algorithm, to reduce computational cost, we fix $\left(\lambda_{1},\lambda_{2}\right)=(0.5\sqrt{n\log p},$ $0.5\sqrt{n\log p})$, $p_{1}=0.85$, and $v_{1}=1$. For the choice of $v_{0}$ and $u$, we adopt the AIC criterion defined as $-2\log f\left(\bm{t},\bm{\delta},\textbf{X}\mid\hat{\bm{\beta}},\hat{\bm{\mu}},\widetilde{\bm{\Omega}};\hat{\bm{\tau}},\hat{\bm{\pi}}_{g}\right)+2\sum_{k=1}^{K}\hat{s}_{k},$ where $\hat{\bm{\beta}},\hat{\bm{\mu}},\widetilde{\bm{\Omega}},\hat{\bm{\tau}}$, and $\hat{\bm{\pi}}_{g}$ are the final updates from the proposed algorithm and $\hat{s}_{k}=\#\left\{(j,l):\widetilde{\omega}_{k,jl}\neq 0,1\leq l<j\leq p\right\}$.
@@ 2.5 Theoretical properties
Furthermore, for a $q$-dimensional vector $\bm{\nu}$, denote $\left\|\bm{\nu}\right\|_{\max}=\max_{1\leq j\leq q}\left|\nu_{j}\right|$, and for a matrix $\bm{M}\in\mathbb{R}^{q_{1}\times q_{2}}$, define its min, max, and max induced norms as $\|\bm{M}\|_{\min}=\min_{i,j}\left|M_{ij}\right|$, $\|\bm{M}\|_{\max}=\max_{i,j}\left|M_{ij}\right|$, and $\|\bm{M}\|_{\infty}=\max_{i=1,\ldots,q_{1}}\sum_{j=1}^{q_{2}}\left|M_{ij}\right|$, respectively. For a square matrix $\bm{A}\in\mathbb{R}^{p\times p}$, denote $\lambda_{\min}(\bm{A})$ and $\lambda_{\max}(\bm{A})$ as its smallest and largest eigenvalues. The following conditions are assumed.
@@
$\underline{C}_{1}\leq\min_{k}\lambda_{\min}\left(\bm{\Omega}_{k}^{*}\right)\leq\max_{k}\lambda_{\max}\left(\bm{\Omega}_{k}^{*}\right)\leq\bar{C}_{1}$, where $\underline{C}_{1}$ and $\bar{C}_{1}$ are positive constants. Moreover, let ¨$\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ and ¨$\ddot{\ell}_{2}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ be the second derivatives of $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}\delta_{i}\log f_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ and $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}(1-\delta_{i})\log S_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ with respect to $\operatorname{vec}\left({\bm{\beta}},\bm{\tau}\right)$, respectively. Then for any $\underline{\bm{\Psi}}\in\Xi$, $\bm{\Psi}\in\mathcal{B}_{\alpha}\left(\bm{\Psi}^{*}\right)=\left\{\bm{\Psi}\in\Xi:\left\|\bm{\Psi}-\bm{\Psi}^{*}\right\|_{2}\right.$ $\left.\leq\alpha\right\}$ and $\bar{\bm{\Psi}}\in\left\{\bar{\bm{\Psi}}\mid\left\|\bar{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}\leq 2\alpha\right\}$, there exists $\widetilde{\eta}>0$ satisfying $\lambda_{\max}\left(\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)\right)\leq-\widetilde{\eta}$ and $\lambda_{\max}\left(\ddot{\ell}_{2}\left(\bm{\Psi},\underline{\bm{\Psi}}\right)\right)\leq 0.$
@@
$\left\|\bm{\beta}^{*}\right\|_{\max}$, $\left\|\bm{\tau}^{\ast}\right\|_{\max}$, $\left\|\bm{\mu}^{*}\right\|_{\max}$, and $\max_{k}\left\|\bm{\Omega}_{k}^{*}\right\|_{\infty}$ are bounded from above, and $\left\|\bm{\tau}^{\ast}\right\|_{\min}$ is bounded from below. Furthermore, $\max\limits_{i=1,\cdots,n}{c}_{i}\leq\bar{C}_{2}$ and $\max\limits_{i=1,\cdots,n}\left\|\bm{x}_{i}\right\|_{\max}\leq\bar{C}_{3}$ for some constants $\bar{C}_{2}$ and $\bar{C}_{3}$.
@@
The tuning parameters satisfy $u=O\left(\sqrt{{n(Ks+p)\log p}/{(Kp^{2}s^{2})}}\right)$, $1/v_{0}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, $1/v_{1}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, and $v_{0}<v_{1}$.
Assumption 1 is a restricted strong concave condition for the log-likelihood function. Assumption 2 is the sufficient separability condition, which requires that the probability of each subject belonging to a subgroup is either close to 1 or 0. Such an assumption is often assumed in the literature [11, 20]. Assumption 3 puts constraints on the bounds of the true parameters and the predictors, which are very common in the relevant literature. Assumption 4 allows $K$ to grow at a rate slower than $\sqrt{p/\log n}$, which is also considered in the related literature [11, 20]. Assumption 5 restricts the rate of the tuning parameters, which are commonly assumed in high-dimensional studies [26, 20, 28].
@@
In Theorem 1, if $K$ is fixed, the non-asymptotic bound of the estimation error reduces to $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\left(d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{{s+p}}\right)\sqrt{\log p/n}\right)$. Denote by $\mathcal{\bm{Q}}\equiv\left\{\bm{\Omega}=\operatorname{vec}\left(\bm{\Omega}_{1},\ldots,\right.\right.$ $\left.\left.\bm{\Omega}_{K}\right):\bm{\Omega}_{k}\text{ is positive definite and~{}}\max_{k}\|\bm{\Omega}_{k}\|_{\infty}\leq C_{\mathcal{Q}}\right\}$ the class of precision matrices. Under Assumption 3, the upper bound $C_{\mathcal{Q}}$ of $\mathcal{\bm{Q}}$ is independent of $(n,p)$. Provided that $d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}=O(\sqrt{s+p})$, the convergence rate of the precision matrix estimate $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}(\sqrt{(s+p)\log p/n})$, which achieves the minimax optimal rate for estimating an $s$-sparse precision matrix under the Frobenius norm as shown in [3].

### v2

@@ 2.4 Posterior Inference
with $f\left(\bm{0}\mid\bm{\theta}\right)=\prod_{j=1}^{p}\prod_{l>j}f\left(\bm{0}\mid\bm{\theta}_{jl}\right)$. We note that since the distribution $f\left(\bm{\theta}\mid\bm{\Omega}\right)$ is the indicator function, there is no need to estimate $\bm{\theta}$ separately. Figure S1 in the supplementary materials gives a graphical representation of the Bayesian model.
Based on (2), we develop an Expectation Maximization (EM) algorithm for posterior inference of the parameter set $\bm{\Psi}=\operatorname{vec}\left(\bm{\beta},\bm{\tau},\bm{\mu},\bm{\Omega}\right)$. Sensitivity analyses on hyperparameters confirm that our approach is robust across a range of plausible settings. The Bayesian Information Criterion (BIC) is adopted to choose optimal values of $v_{0}$ and $u$. We refer to Sections S2 and S3 of the supplementary materials for details.
@@ 2.5 Theoretical properties
Furthermore, for a $q$-dimensional vector $\bm{\nu}$, denote $\left\|\bm{\nu}\right\|_{\max}=\max_{1\leq j\leq q}\left|\nu_{j}\right|$. For a matrix $\bm{M}\in\mathbb{R}^{q_{1}\times q_{2}}$, define its minimum, maximum, and max-induced norms as $\|\bm{M}\|_{\min}=\min_{i,j}\left|M_{ij}\right|$, $\|\bm{M}\|_{\max}=\max_{i,j}\left|M_{ij}\right|$, and $\|\bm{M}\|_{\infty}=\max_{i=1,\ldots,q_{1}}\sum_{j=1}^{q_{2}}\left|M_{ij}\right|$, respectively. For a square matrix $\bm{A}\in\mathbb{R}^{p\times p}$, let $\lambda_{\min}(\bm{A})$ and $\lambda_{\max}(\bm{A})$ denote its smallest and largest eigenvalues. The following conditions are assumed.
@@
Suppose $\underline{C}_{1}\leq\min_{k}\lambda_{\min}\left(\bm{\Omega}_{k}^{*}\right)\leq\max_{k}\lambda_{\max}\left(\bm{\Omega}_{k}^{*}\right)\leq\bar{C}_{1}$, where $\underline{C}_{1}$ and $\bar{C}_{1}$ are positive constants. Moreover, we denote ¨$\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ and ¨$\ddot{\ell}_{2}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ as the second derivatives of $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}\delta_{i}\log f_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ and $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}(1-\delta_{i})\log S_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ with respect to $\operatorname{vec}\left({\bm{\beta}},\bm{\tau}\right)$, respectively. Then for any $\underline{\bm{\Psi}}\in\Xi$, $\bm{\Psi}\in\mathcal{B}_{\alpha}\left(\bm{\Psi}^{*}\right)=\left\{\bm{\Psi}\in\Xi:\left\|\bm{\Psi}-\bm{\Psi}^{*}\right\|_{2}\right.$ $\left.\leq\alpha\right\}$ and $\bar{\bm{\Psi}}\in\left\{\bar{\bm{\Psi}}\mid\left\|\bar{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}\leq 2\alpha\right\}$, there exists $\widetilde{\eta}>0$ satisfying $\lambda_{\max}\left(\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)\right)\leq-\widetilde{\eta}$ and $\lambda_{\max}\left(\ddot{\ell}_{2}\left(\bm{\Psi},\underline{\bm{\Psi}}\right)\right)\leq 0.$
@@
$\left\|\bm{\beta}^{*}\right\|_{\max}$, $\left\|\bm{\tau}^{\ast}\right\|_{\max}$,
$\left\|\bm{\mu}^{*}\right\|_{\max}$, and $\max_{k}\left\|\bm{\Omega}_{k}^{*}\right\|_{\infty}$ are bounded from above, and $\left\|\bm{\tau}^{\ast}\right\|_{\min}$ is bounded from below. Furthermore, $\max\limits_{i=1,\cdots,n}{c}_{i}\leq\bar{C}_{2}$ and $\max\limits_{i=1,\cdots,n}\left\|\bm{x}_{i}\right\|_{\max}\leq\bar{C}_{3}$ for some constants $\bar{C}_{2}$ and $\bar{C}_{3}$.
@@
Assume the tuning parameters satisfy $u=O\left(\sqrt{{n(Ks+p)\log p}/{(Kp^{2}s^{2})}}\right)$, $\tau_{0}=O\left(\sqrt{n{(Ks+p)K\log p}/p}\right)$, $1/v_{0}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, $1/v_{1}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, and $v_{0}<v_{1}$.
Assumption 1 is a restricted strong concave condition for the log-likelihood function [13, 2]. Assumption 2 is the sufficient separability condition, which requires that the probability of each subject belonging to a subgroup is either close to 1 or 0. Such an assumption is often assumed in the literature [12, 22]. Assumption 3 puts constraints on the bounds of the true parameters and the predictors, which are very common in the relevant literature. Assumption 4 allows $K$ to grow at a rate slower than $\sqrt{p/\log n}$, which is also considered in the related literature [17, 8]. Assumption 5 restricts the rate of the tuning parameters, which are commonly assumed in high-dimensional studies [28, 31].
@@
In Theorem 1, if $K$ is fixed, the non-asymptotic bound of the estimation error reduces to $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\left(d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{{s+p}}\right)\sqrt{\log p/n}\right)$. Denote by $\mathcal{\bm{Q}}\equiv\left\{\bm{\Omega}=\operatorname{vec}\left(\bm{\Omega}_{1},\ldots,\right.\right.$ $\left.\left.\bm{\Omega}_{K}\right):\right.$ $\left.\bm{\Omega}_{k}\text{ is positive definite and~{}}\max_{k}\|\bm{\Omega}_{k}\|_{\infty}\leq C_{\mathcal{Q}}\right\}$ the class of precision matrices. Under Assumption 3, the upper bound $C_{\mathcal{Q}}$ of $\mathcal{\bm{Q}}$ is independent of $(n,p)$. Provided that $d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}=O(\sqrt{s+p})$, the convergence rate of the precision matrix estimate $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}(\sqrt{(s+p)\log p/n})$, which achieves the minimax optimal rate for estimating an $s$-sparse precision matrix under the Frobenius norm as shown in [3]. The proof of Theorem 1 is provided in Section S5 of the supplementary materials.